OpenStudy (anonymous):

Write as single logarithims. a.)log x + log y b.) log t^2- log f I'll give medals!!! and Fan!!!

OpenStudy (anonymous):

log a + log b = log a*b USe that property to solve part a

OpenStudy (anonymous):

log a - log b = log (a/b) Use this property to solve part B

OpenStudy (anonymous):

so its log xy

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{\log(a)+\log(b)=\log(a\times b)}\) \(\normalsize\color{green}{\log(a)-\log(b)=\log(a\div b)}\) \(\normalsize\color{red}{\log(a)\div\log(b)=\log_b(a)}\) there are more rules ••••

OpenStudy (anonymous):

thanks @SolomonZelman @samjordon @jdoe0001

OpenStudy (solomonzelman):

Anytime

OpenStudy (jdoe0001):

yw